can you please help me solve this equation, 2x-y+1=0 and x^2-3x-4-y=y^2
Answer by Alan3354(69443)

can you please help me solve this equation, 2x-y+1=0 and x^2-3x-4-y=y^2
-------
2x-y+1=0 --> y = 2x + 1
----
Sub for y in the 2nd equation.